OpenStudy (anonymous):

Please help! What value of j makes the sequence: 1, j, 23, .... arithmetic?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

1+11=12, 12+11=23, etc.

OpenStudy (anonymous):

1+11=12 12+11=23

OpenStudy (anonymous):

\[d=23-j=j-1\]

OpenStudy (anonymous):

d = 11. a1 = 1 a2 = a1 + d = 1 + 11 = 12 a3 = a2 + d = 12 + 11 = 23 So, j = 12.

OpenStudy (anonymous):

a simple formula for that called arthematic mean =a+b/2=1+23/2=24/2=12:)

OpenStudy (anonymous):

thanks y'all! :)